# True Zero-Shot Inference of Dynamical Systems Preserving Long-Term Statistics

### v2

@@ A.2 Datasets
DynaMix is trained on about $0.6$ million simulated time series of length $T=550$ sampled from $34$ different $3d$ DS with cyclic or chaotic attractors, collected in [29]. Time series were standardized to ensure comparable scaling of all data, and Gaussian noise of $5\%$ of the data standard deviation was added to all dimensions for the ground truth. Attractor dynamics were chosen to reflect different types of behavior, see Fig. 10.
@@
Our test set for DS consists of simulated time series of length $10^{5}$ sampled from $54$ different $3d$ DS collected in [29], which are not part of the training set.
Furthermore, we evaluated several $2d$ systems: The Selkov system [77] describing a kinetic model of an open monosubstrate enzyme reaction given by the equations
@@
where we chose $a=0.1$ and $b=0.5$. The Van-der-Pol system [87] describes self-sustaining oscillations in vacuum tubes by
@@
where we chose $\mu=0.5$. Similar to the Van-der-Pol system, the Rayleigh oscillator [61] describes a self-sustained nonlinear oscillator through
@@
where we chose $\mu=1.0$. As a higher dimensional DS test case we use the Lorenz-96 system [54] defined by
@@
We further probed forecasting and DSR on different types of real-world data: The traffic data are hourly recordings of the number of cars passing road junctions (https://www.kaggle.com/datasets/fedesoriano/traffic-prediction-dataset/data). The cloud data, also used in [84] to evaluate TS foundation models, are publicly available from Huawei Cloud (https://github.com/sir-lab/data-release). It consists of function requests from Huawei’s serverless cloud platform. The weather data consists of daily sampled soil temperature and air pressure measured in the city of Mannheim, Germany (source: Deutscher Wetterdienst [German weather service]), and can be accessed via https://www.dwd.de/EN/ourservices/cdc/cdc_ueberblick-klimadaten_en.html. The functional magnetic resonance imaging (fMRI) data comes from human subjects performing cognitive tasks and is publicly available on GitHub [50]. We followed Kramer et al. [50] and selected the first principal component of BOLD activity in each of the $20$ brain regions. Kramer et al. [50] report a positive maximum Lyapunov exponent for models reconstructed from these time series, indicating their chaotic nature (see also [89]). The ETTh1 dataset is part of the Electricity Transformer Temperature (ETT) benchmark, which is widely used for evaluating TS forecasting models. It contains hourly data collected from a power transformer station [100] and can be accessed at https://github.com/zhouhaoyi/ETDataset. Electroencephalogram (EEG) data were taken from a study by Schalk et al. [75], comprising 64-channel data collected from human subjects performing various motor and imagery tasks. Following the approach of Brenner et al. [6], the signals were smoothed using a Hann window of length $15$. As for the training data, all test data were standardized for processing by the models.
@@ A.3 TS foundation models
Chronos is a recent TS foundation framework that adapts a transformer-based LLM architecture for probabilistic time series forecasting [2]. Key to this approach is the tokenization of real-valued time series observations through scaling and quantization, transforming them into token sequences accessible to language models. The model is pretrained on extensive datasets, including both synthetic data generated via Gaussian processes, as well as an extensive batch of empirical time series from [2, 32], including traffic, weather/ climate, electricity and web data, enabling it to achieve strong zero-shot performance across diverse datasets without task-specific fine-tuning. For our evaluation we used the standard pipeline as described in https://github.com/amazon-science/chronos-forecasting.
@@
TimesFM is another transformer based TS foundation model, with a decoder-only style architecture using input patching [20]. Its training corpus consists of synthetic as well as real-world time series data, and it exhibits generalization across different time series domains and context lengths. Evaluation is performed as in https://github.com/google-research/timesfm.
@@
Mamba4Cast is a zero-shot time series forecasting model based on the Mamba architecture [35], a type of linear (‘state-space’) RNN with nonlinear input & output gating, and inspired by Prior-data Fitted Networks (PFNs) [5]. Trained exclusively on synthetic data, Mamba4Cast can generate zero-shot forecasts when provided with time series context information. For evaluation we follow https://github.com/automl/Mamba4Cast.
@@
Like Mamba4Cast, Tiny Time Mixers is a time series foundation model not founded on transformers [26]. It builds upon the TSMixer backbone [25] to which it adds several key innovations such as adaptive patching, diverse resolution sampling, and multi-resolution prefix tuning to improve generalization across datasets with varying temporal resolutions. Evaluation here is performed as in https://github.com/glehet/TTM1.
@@ A.4 Further results
Figure 23: Zero-shot DSR performance across all 54 test set DS for DynaMix and various TS foundation models as in Fig. 4, but for context length $T_{C}=512$. Numerical details in Table LABEL:tab:performance_DSR_CL512.
@@
Figure 26: DSR performance for different teacher forcing intervals $\tau$, illustrating an optimal value is essential for successful training, see [58, 6]. Error bands = STD
@@ A.5 Using DynaMix for non-stationary time series
Similar as in FEDformers [101] and related TS models which use specialized decomposition and filtering blocks for handling non-stationary data, one could add simple preprocessing operations to the DynaMix pipeline to separate out trend or other non-stationary components in the context signal. For a simple illustration, here we first apply a Box-Cox transformation to time series $x_{t}$,
@@ A.6 Scaling up context dimension
For reconstructing DS higher-dimensional than the context dimension defined by the architecture and used in training, one can use DynaMix without any modification by employing the delay-embedding theorems [80, 74]: Zero-shot infer the underlying DS from the observed TS and then delay-embed DynaMix’ output into a sufficiently high-dimensional space which assures a diffeomorphism between original and reconstructed attractor. This idea is illustrated for the $6d$ Lorenz-96 system in Fig. 33. Here, only the first two dimensions of the simulated Lorenz-96 were provided as context, for which DynaMix then produced long-term forecasts. The good geometrical and temporal agreement between delay-embeddings of both the two ground truth and the forcasted TS confirms that DynaMix has correctly inferred the underlying $6d$ DS despite its own $3d$ structure.
@@ A.7 Comparison of DynaMix to custom-trained DSR models
Our first comparison model is the same AL-RNN used for DynaMix’ experts (with $\approx 500$ trainable parameters), i.e. given by eq. 1. Training was performed as in Brenner et al. [7], and hyperparameter settings, as collected in Table 3, followed those in the original paper.
@@
As another SOTA DSR model we trained Neural-ODEs [17], using the same type of MLP architecture, training setup, and same hyper-parameter settings as in [37], see Table 4. With this, the total number of trainable parameters is already more than twice (!) than that used for DynaMix ($\approx 10$k), and exceeded the number of data points ($\leq 4$k) manifold. Indeed, results became worse with less parameters, potentially indicating that Neural-ODEs particularly struggle on the short training segments.
@@
Finally, Reservoir Computers (RCs) are widely used for DSR. Here we employed the architecture provided in Patel and Ott [65], trained as in Göring et al. [37] and using the same hyper-parameter settings, summarized in Table 5. Note that this endows RCs with about $25\times$ more parameters in total, of which $\leq 3$k were trainable, than DynaMix had at its disposal.
@@
Table 9: Performance comparison for zero-shot inference across all DS in the dataset from [29] using $T_{C}=2000$. Best value in red, second best in blue.
System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM TimesFM

### v3

@@ A.2 Datasets
DynaMix is trained on about $0.6$ million simulated time series of length $T=550$ sampled from $34$ different $3d$ DS with cyclic or chaotic attractors, collected in Gilpin [2022]. Time series were standardized to ensure comparable scaling of all data, and Gaussian noise of $5\%$ of the data standard deviation was added to all dimensions for the ground truth. Attractor dynamics were chosen to reflect different types of behavior, see Fig. 10.
@@
Our test set for DS consists of simulated time series of length $10^{5}$ sampled from $54$ different $3d$ DS collected in Gilpin [2022], which are not part of the training set.
Furthermore, we evaluated several $2d$ systems: The Selkov system SEL’KOV [1968] describing a kinetic model of an open monosubstrate enzyme reaction given by the equations
@@
where we chose $a=0.1$ and $b=0.5$. The Van-der-Pol system Van der Pol [1926] describes self-sustaining oscillations in vacuum tubes by
@@
where we chose $\mu=0.5$. Similar to the Van-der-Pol system, the Rayleigh oscillator Nayfeh and Mook [2024] describes a self-sustained nonlinear oscillator through
@@
where we chose $\mu=1.0$. As a higher dimensional DS test case we use the Lorenz-96 system Lorenz [1996] defined by
@@
We further probed forecasting and DSR on different types of real-world data: The traffic data are hourly recordings of the number of cars passing road junctions (https://www.kaggle.com/datasets/fedesoriano/traffic-prediction-dataset/data). The cloud data, also used in Toner et al. [2025] to evaluate TS foundation models, are publicly available from Huawei Cloud (https://github.com/sir-lab/data-release). It consists of function requests from Huawei’s serverless cloud platform. The weather data consists of daily sampled soil temperature and air pressure measured in the city of Mannheim, Germany (source: Deutscher Wetterdienst [German weather service]), and can be accessed via https://www.dwd.de/EN/ourservices/cdc/cdc_ueberblick-klimadaten_en.html. The functional magnetic resonance imaging (fMRI) data comes from human subjects performing cognitive tasks and is publicly available on GitHub Kramer et al. [2022]. We followed Kramer et al. [2022] and selected the first principal component of BOLD activity in each of the $20$ brain regions. Kramer et al. [2022] report a positive maximum Lyapunov exponent for models reconstructed from these time series, indicating their chaotic nature (see also Volkmann et al. [2024]). The ETTh1 dataset is part of the Electricity Transformer Temperature (ETT) benchmark, which is widely used for evaluating TS forecasting models. It contains hourly data collected from a power transformer station Zhou et al. [2021] and can be accessed at https://github.com/zhouhaoyi/ETDataset. Electroencephalogram (EEG) data were taken from a study by Schalk et al. [2004], comprising 64-channel data collected from human subjects performing various motor and imagery tasks. Following the approach of Brenner et al. [2022], the signals were smoothed using a Hann window of length $15$. As for the training data, all test data were standardized for processing by the models.
@@ A.3 TS foundation models
Chronos is a recent TS foundation framework that adapts a transformer-based LLM architecture for probabilistic time series forecasting Ansari et al. [2024]. Key to this approach is the tokenization of real-valued time series observations through scaling and quantization, transforming them into token sequences accessible to language models. The model is pretrained on extensive datasets, including both synthetic data generated via Gaussian processes, as well as an extensive batch of empirical time series from Ansari et al. [2024], Godahewa et al. [2021], including traffic, weather/ climate, electricity and web data, enabling it to achieve strong zero-shot performance across diverse datasets without task-specific fine-tuning. For our evaluation we used the standard pipeline as described in https://github.com/amazon-science/chronos-forecasting.
@@
TimesFM is another transformer based TS foundation model, with a decoder-only style architecture using input patching Das et al. [2024]. Its training corpus consists of synthetic as well as real-world time series data, and it exhibits generalization across different time series domains and context lengths. Evaluation is performed as in https://github.com/google-research/timesfm.
@@
Mamba4Cast is a zero-shot time series forecasting model based on the Mamba architecture Gu and Dao [2023], a type of linear (‘state-space’) RNN with nonlinear input & output gating, and inspired by Prior-data Fitted Networks (PFNs) Bhethanabhotla et al. [2024]. Trained exclusively on synthetic data, Mamba4Cast can generate zero-shot forecasts when provided with time series context information. For evaluation we follow https://github.com/automl/Mamba4Cast.
@@
Like Mamba4Cast, Tiny Time Mixers is a time series foundation model not founded on transformers Ekambaram et al. [2024]. It builds upon the TSMixer backbone Ekambaram et al. [2023] to which it adds several key innovations such as adaptive patching, diverse resolution sampling, and multi-resolution prefix tuning to improve generalization across datasets with varying temporal resolutions. Evaluation here is performed as in https://github.com/glehet/TTM1.
Panda
Panda is a recently proposed foundation model for short-term forecasting of DS Lai et al. [2025]. It is based on a transformer architecture, which relies on patching the DS-generated time series. The model is trained on $2\cdot 10^{4}$ DS produced by combining base DS from the same database used to train and evaluate DynaMix in skew-product form Gilpin [2022]. Evaluation is performed as in https://github.com/abao1999/panda.
@@ A.4 Further results
Figure 23: Zero-shot DSR performance across all 54 test set DS for DynaMix and various TS foundation models as in Fig. 4, but for context length $T_{C}=512$. Numerical details in Table LABEL:tab:performance_DSR_CL512. Statistical testing based on Wilcoxon signed-rank tests.
@@
Figure 26: DSR performance for different teacher forcing intervals $\tau$, illustrating an optimal value is essential for successful training, see Mikhaeil et al. [2022], Brenner et al. [2022]. Error bands = STD
@@ A.5 Using DynaMix for non-stationary time series
Similar as in FEDformers Zhou et al. [2022] and related TS models which use specialized decomposition and filtering blocks for handling non-stationary data, one could add simple preprocessing operations to the DynaMix pipeline to separate out trend or other non-stationary components in the context signal. For a simple illustration, here we first apply a Box-Cox transformation to time series $x_{t}$,
@@ A.6 Scaling up context dimension
For reconstructing DS higher-dimensional than the context dimension defined by the architecture and used in training, one can use DynaMix without any modification by employing the delay-embedding theorems Takens [1981], Sauer et al. [1991]: Zero-shot infer the underlying DS from the observed TS and then delay-embed DynaMix’ output into a sufficiently high-dimensional space which assures a diffeomorphism between original and reconstructed attractor. This idea is illustrated for the $6d$ Lorenz-96 system in Fig. 33. Here, only the first two dimensions of the simulated Lorenz-96 were provided as context, for which DynaMix then produced long-term forecasts. The good geometrical and temporal agreement between delay-embeddings of both the two ground truth and the forcasted TS confirms that DynaMix has correctly inferred the underlying $6d$ DS despite its own $3d$ structure.
@@ A.7 Comparison of DynaMix to custom-trained DSR models
Our first comparison model is the same AL-RNN used for DynaMix’ experts (with $\approx 500$ trainable parameters), i.e. given by eq. 1. Training was performed as in Brenner et al. [2024a], and hyperparameter settings, as collected in Table 3, followed those in the original paper.
@@
As another SOTA DSR model we trained Neural-ODEs Chen et al. [2018], using the same type of MLP architecture, training setup, and same hyper-parameter settings as in Göring et al. [2024a], see Table 4. With this, the total number of trainable parameters is already more than twice (!) than that used for DynaMix ($\approx 10$k), and exceeded the number of data points ($\leq 4$k) manifold. Indeed, results became worse with less parameters, potentially indicating that Neural-ODEs particularly struggle on the short training segments.
@@
Finally, Reservoir Computers (RCs) are widely used for DSR. Here we employed the architecture provided in Patel and Ott [2022], trained as in Göring et al. [2024a] and using the same hyper-parameter settings, summarized in Table 5. Note that this endows RCs with about $25\times$ more parameters in total, of which $\leq 3$k were trainable, than DynaMix had at its disposal.
@@
Table 9: Performance comparison for zero-shot inference across all DS in the dataset from Gilpin [2022] using $T_{C}=2000$. Best value in red, second best in blue.
System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-2 Panda Mamba4Cast TTM TimesFM